derive;

Wnc=(½mvf2+mghf)-(½mv02+mgh0)


1
Expert's answer
2021-11-08T09:12:37-0500

For kinetic energy 


"m\frac{dv_x}{dt}=F_x\to\frac{d}{dt}(\frac{mv_x^2}{2})=F_xv_x \to d(\frac{mv_x^2}{2})=F_xdx\to"


"\frac{mv_{xf}^2}{2}-\frac{mv_{x0}^2}{2}=\int F_xdx" . In general "\frac{mv_{f}^2}{2}-\frac{mv_{0}^2}{2}=\int \vec Fd\vec r"


For potential energy


"\vec Fd\vec r=-\frac{\partial E_p}{\partial x}dx-\frac{\partial E_p}{\partial y}dy-\frac{\partial E_p}{\partial z}dz \to \vec Fd\vec r=-dE_p \to"


"-(E_{p2}-E_{p1})=\int \vec Fd\vec r"


If the body has both potential and kinetic energy then


"W=\int \vec Fd\vec r=\frac{mv_{f}^2}{2}-\frac{mv_{0}^2}{2}-(-(E_{p2}-E_{p1}))="


"=(\frac{mv_{f}^2}{2}+E_{p2})-(\frac{mv_{0}^2}{2}+E_{p1})" .